OpenStudy (anonymous):

Find the zeros (real and/or imaginary). f(x)=x^2+9

zepdrix (zepdrix):

Ooo it's Hello Kitty c:

OpenStudy (anonymous):

set it equal to zero and solve for x

OpenStudy (anonymous):

Haha zepdrix :)

OpenStudy (anonymous):

so, \[0 = x^2 +9\]\[-9 = x^2\]\[\sqrt{-9}=x\]\[\pm3i = x\]

OpenStudy (anonymous):

wait but I have two x's

OpenStudy (anonymous):

oh wait nevermind, I was looking at a different problem. okay thanks!

OpenStudy (anonymous):

np